Problem Page Two buses leave towns 865km apart at the same time and travel toward each other. One bus travels 17 /kmh slower than the other. If they meet in 5 hours, what is the rate of each bus?
Answer by stanbon(75887)

Problem Page Two buses leave towns 865km apart at the same time and travel toward each other. One bus travels 17 /kmh slower than the other. If they meet in 5 hours, what is the rate of each bus?
-----
slower bus DATA:
rate = r km/hr ; time = 5 hrs ; dist = 5r km
--------------------
faster bus DATA:
rate = r+17 km/hr ; time = 5 hrs ; dist = 5(r+17) km
----------------------
5r + 5(r+17) = 865
10r + 85 = 865
10r = 780
rate = 78 km/hr (rate of slower car)
---
r+17 = 93 km/hr (rate of faster car)
-------------
Cheers,
Stan H.
-----------